OpenStudy (mtalhahassan2):

Need help!! Solve a) log base x 625 =4

geerky42 (geerky42):

Hint: \(\log_x625 = \dfrac{\log 625}{\log x}\)

geerky42 (geerky42):

Can you handle it now?

OpenStudy (loser66):

We have equivalent formulas. Those are \(log_a b= c\iff a^c =b\)

OpenStudy (loser66):

like \(log_2 8 = 3 \iff 2^3 =8\) Same as your problem, you can convert from log to exponent and solve for x

OpenStudy (mtalhahassan2):

thnx a lot both of you

OpenStudy (mtalhahassan2):

wait but how can we do this one

OpenStudy (mtalhahassan2):



OpenStudy (mtalhahassan2):